A 8m thick layer of medium sand overlies a deep deposit of dense gravel. A series of standard penetration tests carried out through the depth of the sand has established that the average blow count is 26. Further tests show that the gravel has a standard penetration value of N = 52 in the region of the interface with the sand. A precast pile of diameter 300mm is to be driven down through the sand and to penetrate sufficiently into the gravel to give good end bearing. Adopting a safety factor of 3.0, determine the allowable load that the pile will be able to carry.



1
Expert's answer
2021-05-24T17:25:01-0400

given that,

Average blow count "=26"

diameter of forecast file "=300mm=0.3m"

standard penetration value (N)"=52"

factor of safety(fs)"=3"

length of file "=8m=8000mm"

allowable load"=?"

Qu"=400NA_b+2\bar{N}A_s"

where,

"A_b-" gross sectional area

i.e "A_b=\frac{\pi}{4}\times d^
2"

"=\frac{\pi}{4}(300)^2"

"=70685.8mm^2" (Q)"0.07068"

"A_s=" lateral surface area

i.e, "A_s=\pi\times" dia of file "\times" length of file

"=\pi \times d\times l"

"=\pi\times300\times8000"

"=\pi\times2400000"

"=7539822.37mm^2" (Q) "7.54m^2"

Qu"=(400)(52)(0.0707)+(2)(26)(7.54)"

Qu"=14705.6+392.08"

allowabe load"=\frac{Qu}{0.5}"

"=1509.77\/3"

"=502.25KN"